For which values of a and b is the following equation true?
lim((sin(2x)/x^3)+a+(b/x^2))=0,x approaches to 0
1
Expert's answer
2020-03-19T11:57:43-0400

The limit equation is given by:

"\lim_{x \to 0} \left( \frac{sin(2x)}{x^3} + a + \frac{b}{x^2} \right) = 0"

Separating the constant "a" from the limit, we have:

"\lim_{x \to 0} \left( \frac{\sin (2x)}{x^3} + \frac{b}{x^2} \right) + a = 0"

"\Rightarrow \lim_{x \to 0} \left( \frac{ \sin (2x) + bx }{x^3} \right) + a = 0"

After substituting the limit value "x = 0" into the limit function, we can see that we have "\frac{0}{0}" form is exist. So we will apply L'Hospital's Rule to this limit function.

Now:

We know that this rule states that if such indeterminate form "\frac{0}{0}" exists then we will take the derivatives of the numerator and denominator and substitute the limit value. This process will apply again and again until we get a finite value.

"\\"

Applying L'Hospital's Rule to the limit function, we have:

"\lim_{x \to 0} \left( \frac{2 \cos (2x) + b }{3x^2} \right) + a = 0"

Substituting "x = 0" into the limit function, we have:

"\Rightarrow \left( \frac{2 + b}{3(0^2)} \right) + a \hspace{1 cm} \left[ \because \cos (0) = 1 \right]"

After substituting "x = 0" into the limit function, we get "\infty" . But this result is not possible because the right hand side of the limit equation is zero. So, we have to find a value of "b" so that "b + 2 = 0 \, or \, b = - 2"

"\\"

Now:

Substituting "b = - 2" into the limit function we have:

"\lim_{x \to 0} \left( \frac{2 \cos(2x) - 2}{3x^2} \right) + a = 0"

"\Rightarrow \frac{2}{3} \lim_{x \to 0} \left( \frac{\cos (2x) - 1}{x^2} \right) + a = 0"

"\Rightarrow \frac{2}{3} \lim_{x \to 0} \left( \frac{ - 2 \sin (2x) - 0}{2x} \right) + a = 0" "\hspace{1 cm}" [ "\because \frac{0}{0}" form, so we used L'Hospital's Rule ]

"\Rightarrow - \frac{2}{3} \lim_{x \to 0} \left( \frac{ \sin (2x)}{x} \right) + a = 0"

"\Rightarrow - \frac{2}{3} \lim_{x \to 0} \left( \frac{2 \cos(2x)}{1} \right) + a = 0 \hspace{1 cm}" [ "\because \frac{0}{0}" form, so we used L'Hospital's Rule ]

"\Rightarrow - \frac{2}{3} \left( 2 \cos(0) \right) + a = 0 \hspace{1 cm}" [ Substitute "x = 0" into the limit function ]

"\Rightarrow - \frac{4}{3} + a = 0 \hspace{1 cm} \left[ \because \cos(0) = 1 \right]"

"\Rightarrow a = \frac{4}{3}"

Hence:

The values of "a\, \, and \,\, b" are "\frac{4}{3} \, and \, -2" respectively.